find a G.P for which the sum of the first two terms is -4 and fifth term is 4 times the third term.

Grade:8

1 Answers

Raadhesh Kannan
8 years ago
a+ar=-4 ----(1)
ar^4= 4 * ar^2------(2)
from (2), r=2 ----(3)
 
Using (3) in (1), we get a = -4/3.
 
….….….….….….….….….….….….….….….….….END….….….….….….….….….….….….….….….….….….…..